OpenStudy (help108):

MEDAL AND FAN!!!! What is the solution to the equation fraction 3 over 5 n minus fraction 4 over 5 equals fraction 1 over 5 n?

OpenStudy (help108):

@AloneS

OpenStudy (help108):

@iGreenHacked

OpenStudy (help108):

@AloneS

OpenStudy (anonymous):

\[3/5n - 4/5 = 1/5n\] Flip everything \[5n/3-5/4=5n\] Subtract 5n/3 from both sides \[-5/4=5n-5n/3\] Multiply the "5n" by 3 so that you can subtract \[-5/4=15n/3 - 5n/3\] Subtract \[-5/4=10n/3\] Multiply both sides by 3 \[-15/4=10n\] Divide both sides by 10 \[n=-15/40\] simplify \[n=-3/8\]

OpenStudy (help108):

ok

OpenStudy (help108):

What is the solution to the equation fraction 3 over 5 n minus fraction 4 over 5 equals fraction 1 over 5 n? n = −1 n = 2 n = fraction 5 over 3 n = fraction 5 over 2

OpenStudy (help108):

which option is it B or A?

OpenStudy (anonymous):

n=1/2\[\text{Expand}\left[5 n \left(\frac{3}{5 n}-\frac{4}{5}\right)=\frac{5 n}{5 n}\right] \]\[3-4 n=1 \]\[4 n=2 \]\[n\to \frac{1}{2} \]

OpenStudy (anonymous):

whoops, robtobey is right but that isn't an answer?

OpenStudy (anonymous):

n = 1/2\[\left\{\frac{3-4 n}{5 n},\frac{1}{5 n}\right\}\text{/.}\, n\to \frac{1}{2} \]\[\left\{\frac{2}{5},\frac{2}{5}\right\} \]

OpenStudy (anonymous):

-> is a Mathematica operator that means the LHS can be replaced by the RHS .

OpenStudy (help108):

so which is it a b c or d

OpenStudy (help108):

@robtobey

OpenStudy (help108):

HELPPPPPP!!!!!!!

OpenStudy (help108):

@AloneS

OpenStudy (help108):

which option plzzz help......

OpenStudy (help108):

this is my lifeeeeeeeeeeeeeeeeee

OpenStudy (help108):

help please anyone alone plzzzzzz

OpenStudy (help108):

@Elsa213

alones (alones):

@robtobey already explained it x3

OpenStudy (help108):

which option is it? a b c or d

OpenStudy (help108):

i medaled u AloneS but which one is it?

OpenStudy (help108):

What is the solution to the equation fraction 3 over 5 n minus fraction 4 over 5 equals fraction 1 over 5 n? n = −1 n = 2 n = fraction 5 over 3 n = fraction 5 over 2

OpenStudy (help108):

@AloneS

OpenStudy (help108):

@AloneS